# Towards AI Accountability Infrastructure: Gaps and Opportunities in AI Audit Tooling

### v1

@@ Abstract.
Audits are critical mechanisms for identifying the risks and limitations of deployed artificial intelligence (AI) systems. However, the effective execution of AI audits remains incredibly difficult. As a result, practitioners make use of various tools to support their efforts. Drawing on interviews with 35 AI audit practitioners and a landscape analysis of 390 tools, we map the current ecosystem of available AI audit tools. While there are many tools designed to assist practitioners with setting standards and evaluating AI systems, these tools often fell short of supporting the accountability goals of AI auditing in practice. We thus highlight areas for future tool development beyond evaluation—from harms discovery to advocacy—and outline challenges practitioners faced in their efforts to use AI audit tools. We conclude that resources are lacking to adequately support the full scope of needs for many AI audit practitioners and recommend that the field move beyond tools for just evaluation, towards more comprehensive infrastructure for AI accountability.
@@ 1. Introduction
A common thread weaving through most AI policy proposals—from the E.U. Digital Services Act to a municipal hiring bill passed in New York City (cumbo_local_2021, 26)—is the now-familiar call for AI audits. Often defined as independent evaluations of the performance, fairness, or safety of deployed AI systems, AI audits have now been featured in several U.S. congressional bills (clarke_algorithmic_2019, 23, 59) and state efforts (perrigo_california_2023, 74, 66), and the practice is regularly mentioned in AI governance proposals internationally (galindo_overview_2021, 39).
@@
Despite this increasing policy enthusiasm, the execution of effective AI audits remains practically difficult. The maturity of the audit ecosystem in the technology sector still lags far behind audit practices common in other industries, such as finance and healthcare (raji_closing_2020, 80, 81). A dearth of generalized AI audit guidance is remedied only partially by recent efforts from government advisory bodies like the U.K.’s Information Commissioner’s Office (ICO) (information_commissioners_office_annex_2023, 49), the U.S. National Institute of Standards and Technology (tabassi_ai_2023, 99) and others (office_of_science_and_technology_policy_blueprint_2022, 72). In actuality, these evaluations remain inconsistent and unreliable (ryan-mosley_why_2023, 83), and the lack of access and visibility to many AI products leaves external auditors without the information needed to make adequate, truly legitimate independent assessments (holstein_improving_2019, 48).
In the face of these challenges, practitioners often rely on tools—software, frameworks, and other resources—to support their AI audit work. Past research in human-computer interaction (HCI) and social computing has developed and studied a host of fairness, explainability, and other toolkits that inform such evaluations (bellamy_ai_2018, 9, 95, 51, 12, 29, 112, 17, 65, 32, 58, 48, 30, 5, 111, 35), and governments across the world are developing similar tools for AI governance (kaye_risky_2023, 52). Studies of these evaluation toolkits are crucial for ensuring that AI systems are evaluated thoroughly and that stakeholders are included in the process. However, evaluation alone is not sufficient to hold AI builders and operators responsible and answerable for system behavior (goodman_ai_2022, 44, 81)—auditors also need tools that support key components of accountability such as auditor independence, data access, peer review, standardization, and advocacy (costanza-chock_who_2022, 24).
@@
Figure 1. Stages of the tool-supported audit process surfaced in our survey of AI audit tooling. We taxonomize tools by the stage of the AI audit process in which they are meant to be used.
@@
Figure 1. Stages of the tool-supported audit process surfaced in our survey of AI audit tooling. We taxonomize tools by the stage of the AI audit process in which they are meant to be used.
In this study, we survey the landscape of AI audit tooling to understand the challenges to accountability and potential opportunities for tool development and research. We taxonomized 390 audit tools and interviewed 35 audit practitioners employed at 24 different tech companies, startups, government agencies, non-profits, consulting firms, and academic institutions to explore the following questions:
  1. (1)
    RQ1: What AI audit tooling is available to support AI audit work?
  2. (2)
    RQ2: Which AI audit tools are actually used and how?
  3. (3)
    RQ3: What do AI audit practitioners actually need?
To investigate RQ1, we identified a landscape of audit tools that maps into seven key stages of tool-supported AI auditing—including both tools for evaluation and tools for other important aspects of auditing, such as harms discovery or advocacy. To investigate RQ23, we interviewed AI audit practitioners directly about their experience with AI audit tooling and which gaps they could identify.
We find that while there there are many tools to support audit work, particularly for evaluating AI systems and managing standards, these tools often fell short of helping auditors achieve accountability in practice. Some tools were empowering—but tools to help with tasks outside evaluation, such as discovering harms, communicating audit results, and advocating for subsequent changes, were much less common. Instead, the practitioners we interviewed often adapted existing tools or built their own from scratch to relieve the tedious or difficult tasks in their particular audit workflows. Auditors struggled to find tools, obtain high-quality, uncompromised data, apply consistent and holistic standards and methods, and ensure audit integrity—key challenges in ensuring that audits are effective and impactful.
Our results identify challenges in every stage of tool-supported AI auditing, but we also include suggestions for researchers, policymakers, and practitioners that may provide an alternative vision for tool development that supports accountability, inclusion, and independence. We conclude with a summary of broad lessons for the HCI community and other stakeholders that could help push the landscape of AI audit tools beyond evaluation and towards infrastructure for meaningful accountability.
@@ 2. Related work
AI audits involve the independent evaluation of the performance, risk and safety of deployed AI systems (power_audit_1999, 76, 81). These evaluations are required in order to make consequential judgements about the suitability of an AI product for widespread or context-specific use, and assess how well deployed systems match stakeholder claims and expectations. In this case we use the term artificial intelligence (AI) broadly, to refer to any AI-advertised product or model, including automated decision systems (ADS), algorithmic recommendation systems, large machine learning base models and more.
Past work on AI audit tools. Recent research in human-computer interaction (HCI), social computing, and cooperative design documents the experiences of practitioners evaluating AI and the challenges they face (costanza-chock_who_2022, 24, 48, 17, 65, 32, 58). For example, (holstein_improving_2019, 48, ) documents the practical and technical difficulties faced by internal auditors of ML systems trying to identify and improve fairness, while other studies have examined the organizational challenges and barriers these practitioners face (rakova_where_2021, 82, 65, 109, 86, 24). Several HCI studies have since specifically examined the ways AI audit practitioners make use of various tools to address these issues (deng_exploring_2022, 30, 58, 48, 5). In this study, we extend these analyses to lay out a more complete view on the utility of such tools, which we designate as AI audit tools: any software, services, frameworks, and other artifacts used to conduct audits of AI systems.
While some studies of AI audit tools focus on performance analysis (amershi_modeltracker_2015, 5) or user-driven grassroots auditing (devos_toward_2022, 32, 29), most HCI studies focus specifically on tools for assessing fairness. (lee_landscape_2021, 58, ), for example, compare six prominent open source fairness toolkits along various criteria related to practitioners’ needs. (deng_exploring_2022, 30, ) documented the ways practitioners learn about and use two prominent fairness toolkits, AI Fairness 360 and Fairlearn. And a survey of 152 audit practitioners, (costanza-chock_who_2022, 24, ) found that 62% of practitioners used existing tools like AI Fairness 360, Scikit Fairness, or Parity, though only 7% of respondents used a standardized framework for their overall audit protocol.
Accountable auditing. Our study aims to expand HCI research by looking beyond fairness toolkits to examine the broader array of AI audit tools available to practitioners. Evaluation represents only one phase in a broader process. In addition to evaluation, an effective audit both identifies harms and formulates effective strategies for either addressing them or for holding relevant stakeholders accountable111 We use the term “accountability” in the legal-political sense, meaning to face consequential judgement for a systems’ behaviors and impacts (bovens_analysing_2007, 16, 31). for the audit result (raji_actionable_2022, 78, 80, 24).
@@
The practice of AI auditing, or algorithm auditing—a term first formally proposed by (sandvig_auditing_2014, 84, ) to describe methods for detecting discrimination in online platforms—has expanded over the last decade (vecchione_algorithmic_2021, 104). Researchers have since used the term to encompass not just field studies for detecting discrimination with causal estimation(metaxa_auditing_2021, 67) but also any kind of independent assessment of an automated or data-defined system (digital_regulation_cooperation_forum_auditing_2022, 33, 80). Early audit studies—of facial recognition systems and criminal risk assessment (angwin_machine_2016, 6, 18), for example—by external journalists, academics, and civil advocates fueled advocacy that resulted in advocacy and, at times, even changes to the deployed systems (raji_actionable_2022, 78, 91, 96, 88). However, not all audits succeed in holding system builders and operators accountable (goodman_ai_2022, 44, 108). While these key studies define the practice of AI auditing and some of the components of accountability, no study yet examines the full range of tools used to support AI auditing and accountability. Most notably, (wong_seeing_2023, 111, ) examine documentation from 27 toolkits for “AI Ethics”, finding that these resources employ a narrow technical framing that fails to involve more diverse stakeholders or reckon with the non-technical dimensions of AI ethics work, (kaye_risky_2023, 52, ) survey and critique the tools currently used for AI governance by governments across the globe, and (berman_scoping_2024, 11, ) call for more evaluations of the effectiveness of these tools. We build on these critical studies with a survey of the broader landscape of audit tooling, including some non-technical resources and several previously overlooked areas of tooling necessary for accountability.
@@
Internal vs. external audits. More recently, corporations and policymakers have focused on what we refer to as internal AI audits [^2] Here, the auditors are still considered independent, as they are separate from the development process (i.e., product and engineering teams), even though they remain within the scrutinized organization. (shevlane_model_2023, 92, 36). These audits—afforded better access and cooperation than external audits—may enable accountability if designed properly (raji_closing_2020, 80, 31), but they may also result in false assurances (“audit washing”) (raji_outsider_2022, 81, 44), or foreclose on key remedies such as abolition or disgorgement (sloane_silicon_2022, 94, 61). As a result, there is an equally important role for external AI audits, investigations conducted by civil society, journalists, lawyers, regulators, and other third-party actors. These external audits are typically voluntary research studies and investigations into the deployed AI systems. Past HCI work has not primarily focused on the experiences and challenges of those doing external audit work, and so we have few references for understanding their tooling and resource needs (bandy_problematic_2021, 7).
@@ 3. Methodology
To understand the tools available to audit practitioners, we collated a dataset of 390 tools used or intended for AI audit work and developed a taxonomy of the audit tool landscape based on our findings. To understand exactly which tools auditors use in their everyday work—and where those tools fall short—we also conducted 27 semi-structured interviews with 35 auditors across 24 organizations employing internal and external AI auditors.
@@ 3.1. Interview methodology
To explore the ways audit practitioners are using tools—and what challenges they face—we conducted 27 interviews with a total of 35 audit tool builders and practitioners, representing diverse backgrounds such as engineering, law, journalism, advocacy, policy, and academia in North America ($N=22$) and Europe ($N=5$) (Table 1). Participants for the interviews were recruited with convenience sampling and snowball sampling—we began by contacting practitioners in our professional networks who had conducted notable AI audit work and were active in AI audit communities. Occasionally, participants were referred us to by a colleague or professional contact at another organization. Our sample encompassed both internal and external auditors employed by for-profit tech companies, AI startups, research & civil society non-profits, universities, and government agencies.
@@
Employer
Roles of Interviewees
Participants
Large tech for-profit
Director of Policy Research, VP of Research, Data Science Mgr., Research Eng., Researcher
P5, P8, P11, P14, P19, P27
Tech startup
Co-Founder, CEO, Chief Scientist
P4, P12
Government agency
Tech Policy Principal/Mgr./Assoc./Advisor, Research Fellow
(P21, P28-35), P25
University
Assoc./Asst. Professor, Postdoc. Fellow, Data Scientist
P3, P10, P13, P17, P20
Research non-profit
Co-Founder, Director, Research Scientist
P9, P16, P18
Civil society non-profit
Director, Head of Analytics, Statistician, Researcher, Policy Fellow
P1, P2, P6, P22, P24
Non-profit news org.
Opinion Writer, Data Journalist
P7, P15
Law/consulting for-profit
Policy Director, Mgr., Consultant
(P23, P28), P27
Interviews followed a semi-structured format and lasted 30–60 minutes. Our questions centered on the specific tools and methods practitioners built or employed and 2) common obstacles and unmet needs encountered in their development and use. (The full interview protocol is included in Appendix B.) Participants had the option to remain anonymous and skip questions at their discretion, though none skipped a question. Our protocol was approved by three university IRBs. To analyze the interview data, we transcribed each interview and annotated the transcripts with manual codes. Our coding approach followed an inductive methodology, allowing patterns and themes to emerge from the data (wolcott_transforming_1994, 110, 43). We employed a combination of descriptive coding, which captured the content of the interviews, and values coding, which captured the attitudes and beliefs expressed by participants. Through collaborative sessions and memo writing, we organized these codes and related quotes into key insights, presented in §4.

### v3

@@ Abstract.
Audits are critical mechanisms for identifying the risks and limitations of deployed artificial intelligence (AI) systems. However, the effective execution of AI audits remains incredibly difficult, and practitioners often need to make use of various tools to support their efforts. Drawing on interviews with 35 AI audit practitioners and a landscape analysis of 435 tools, we compare the current ecosystem of AI audit tooling to practitioner needs. While many tools are designed to help set standards and evaluate AI systems, they often fall short in supporting accountability. We outline challenges practitioners faced in their efforts to use AI audit tools and highlight areas for future tool development beyond evaluation—from harms discovery to advocacy. We conclude that the available resources do not currently support the full scope of AI audit practitioners’ needs and recommend that the field move beyond tools for just evaluation and towards more comprehensive infrastructure for AI accountability.
@@ 1. Introduction
Despite increasing policy enthusiasm, [^1] AI audits have been featured in several recent U.S. congressional bills (clarke_algorithmic_2019; lenhart_federal_2023) and state efforts (perrigo_california_2023; mendelson_stop_2021), and the practice is regularly mentioned in AI governance proposals internationally (galindo_overview_2021), from the E.U. Digital Services Act to a municipal hiring bill passed in New York City (cumbo_local_2021). the execution of effective AI audits remains practically difficult. Often defined as independent evaluations of the performance, fairness, legality, or safety of deployed AI systems, the maturity of the audit ecosystem in the technology sector lags far behind other industries such as finance and healthcare (raji_closing_2020; raji_outsider_2022). [^2] This dearth of generalized AI audit guidance is remedied only partially by recent efforts from government advisory bodies like the U.K.’s Information Commissioner’s Office (ICO) (information_commissioners_office_annex_2023), the U.S. National Institute of Standards and Technology (tabassi_ai_2023) and others (office_of_science_and_technology_policy_blueprint_2022). AI audits are often inconsistent and unreliable (ryan-mosley_why_2023), and the lack of access and visibility to many AI systems leaves auditors without the information needed to make adequate and truly independent assessments (holstein_improving_2019; terzis_law_2024).
In the face of these challenges, practitioners often rely on tools—software, frameworks, and other resources—to support their AI audit work. Past research in human-computer interaction (HCI) and social computing has developed and studied a host of fairness, explainability, and other toolkits that inform such evaluations (bellamy_ai_2018; smith-renner_no_2020; kaur_interpreting_2020; bertrand_selective_2023; deng_understanding_2023; woodruff_qualitative_2018; brown_toward_2019; madaio_co-designing_2020; devos_toward_2022; lee_landscape_2021; holstein_improving_2019; deng_exploring_2022; amershi_modeltracker_2015; wong_seeing_2023; ehsan_human-centered_2020). Governments across the globe are developing their own tools and making use of existing resources as part of their enforcement regimes for AI governance (kaye_risky_2023). [^3] Examples include the U.S. AI Safety Institute’s Inspect, the U.S. National Institute of Standards and Technology’s ARIA & Dioptra, Singapore’s AI Verify, and the U.S. National Science Foundation’s Artificial Intelligence Research Resource (NAIRR) Pilot. In the U.S. alone, several recently proposed “AI innovation” bills are explicitly geared towards investing in tooling and resource development for AI auditing. [^4] Examples include the “CREATE AI Act of 2023”, “VET Artificial Intelligence Act”, “Promoting United States Leadership in Standards Act of 2024”, “TEST AI Act of 2023”, “Artificial Intelligence Research, Innovation, and Accountability Act of 2023”, “Artificial Intelligence Public Awareness and Education Campaign Act”, and “Future of Artificial Intelligence Innovation Act of 2024”. Similarly, in the E.U., enforcement reports for the Digital Services Act emphasize the importance of AI audit tooling for effective oversight enforcement (klinger_delegated_2023).
Despite these developments, research has yet to properly map, taxonomize, and understand the full scope of tooling needed to meaningfully support AI audit practitioners. HCI research has identified many practical challenges facing practitioners (holstein_improving_2019; costanza-chock_who_2022; brown_toward_2019; madaio_co-designing_2020; devos_toward_2022) and valuable recent work critically examines some of the toolkits involved (lee_landscape_2021; deng_exploring_2022; wong_seeing_2023; berman_scoping_2024). However, the auditing process involves more than just a performance analysis of the AI product or model. A thorough evaluation alone is not sufficient to hold key stakeholders responsible for system-wide behavior (goodman_ai_2022; raji_outsider_2022)—auditors also need tools that support key components of the accountability process, including target identification, standardization of practice, communication, and advocacy (costanza-chock_who_2022).
In this study, we compare audit practitioners’ tooling needs to the current landscape of AI audit tooling to understand challenges to accountability and potential opportunities for expanding the scope of HCI research and tool development.
  1. (1)
    RQ1: What tools do practitioners need to conduct AI audits?
  2. (2)
    RQ2: What tools currently exist to support AI audit work?
  3. (3)
    RQ3: Do existing tools support practitioners’ needs?
To investigate RQ1, we interviewed 35 AI audit practitioners—employed at 24 organizations, including tech companies, startups, government agencies, non-profits, consulting firms, and academic institutions—about the tools they use and how they support the audit process. To investigate RQ2, we conducted a landscape analysis of 435 existing audit tools (Fig. 1). [^5] Note that our sample is not an exhaustive list of all AI audit tools. In the interviews, we asked practitioners about gaps they encountered and compared their responses to the current tooling landscape (RQ3).
@@
Figure 1. Stages of the tool-supported audit process surfaced in our survey of AI audit tooling. We taxonomize tools by the stage of the AI audit process in which they are used. Tools may be used in multiple stages.
[…]
@@
We find that while there currently exist many tools to support audit work, particularly for evaluating AI systems and managing standards, these tools often fell short of helping auditors achieve accountability in practice. Practitioners found some tools—such as open source tools for data collection—empowering, but tools for tasks beyond evaluation, such as discovering harms, communicating audit results, and advocating for subsequent changes, were much less common. The practitioners we interviewed often adapted existing tools or built their own from scratch to relieve the tedious or difficult tasks in their particular audit workflows. Auditors envisioned tools to help them access high-quality, uncompromised data, apply consistent and holistic standards and methods, and ensure audit integrity.
Our results identify challenges for every stage of the AI audit process as well as opportunities for HCI researchers, policymakers, and practitioners to pursue an alternative vision for tool development that supports rigor, inclusion, independence, and accountability. We conclude with a summary of research and design opportunities for the HCI community and other stakeholders that could help push the landscape of AI audit tools beyond evaluation and towards infrastructure for meaningful accountability.
@@ 2. Related work
AI auditing. The practice of AI auditing, or algorithm auditing—a term first formally proposed in 2014 by sandvig_auditing_2014 to describe methods for detecting discrimination in online platforms—has expanded over the last decade (vecchione_algorithmic_2021). Researchers have since used the term to encompass not just field studies for detecting discrimination with causal estimation (metaxa_auditing_2021) but also any kind of independent assessment of an automated or data-defined system (digital_regulation_cooperation_forum_auditing_2022; raji_closing_2020). Early audit studies of facial recognition systems and criminal risk assessments, (angwin_machine_2016; buolamwini_gender_2018), for example, resulted in widespread advocacy and, at times, even changes to or recalls of the audited systems (raji_actionable_2022; sherwin_facebook_2019; spinks_contemporary_2020; sheard_banning_2021). However, not all audits succeed in holding system builders and operators accountable (goodman_ai_2022; watkins_governing_2021; birhane_ai_2024).
Following birhane_ai_2024, we use the term AI audit (Def. B.2) to refer to “any independent assessment of an identified audit target via an evaluation of articulated expectations with the implicit or explicit objective of accountability”. Accountability is used here in the legal-political sense to mean consequential judgment of a systems’ behaviors and downstream impacts (bovens_analysing_2007). Consequential judgment distinguishes an audit from a simple evaluation or assessment (Fig. 1).
@@
More recently, corporations and policymakers have focused on what we refer to as internal AI audits (Def. B.3) conducted by teams of employees or contractors separate from the product and engineering teams (european_parliament_regulation_2016; raji_outsider_2022). These internal audits, afforded access in cooperation with audit targets, may enable accountability if designed properly (raji_closing_2020; desai_trust_2017), but they may also result in false assurances (“audit washing”) (raji_outsider_2022; goodman_ai_2022) or foreclose on key remedies such as abandonment or disgorgement (Def. B.5) (sloane_silicon_2022; li_algorithmic_2022; johnson_fall_2024). As a result, there is an important role for external AI audits (Def. B.3): investigations conducted by civil society, journalists, lawyers, regulators, and other third-party actors. These external audits are typically voluntary research studies and investigations into deployed AI systems. However, as our findings reveal, external auditors faced significant hurdles to accountability, including lack of access to the systems they aimed to evaluate. The experiences and challenges of those doing external audit work has not been the primary focus of HCI research. Our work provides references for understanding their tooling and resource needs (bandy_problematic_2021; birhane_ai_2024; costanza-chock_who_2022).
While these prior studies define the practice of AI auditing and some of the components of accountability, no study yet examines the full range of tools and technical infrastructure used to support AI auditing and accountability.
Infrastructure & accountability. A rich field of literature in science, technology and society (STS) describes infrastructure as not merely physical or technological, but also as holding social and relational power (winner_artifacts_1980; star_ethnography_1999). Infrastructure is “something that other things ‘run on”’ (lampland_standards_2009): the invisible and axiomatic basis of tools, standards and frameworks that uphold and shape more complex existing processes, services, and engineered artifacts. Embedded in social systems, technological infrastructures encode standards, norms, and guidelines for social organization (verbeek_materializing_2006; lampland_standards_2009).
Accountability also requires infrastructure. Across various industries, the common accountability practice of auditing has long relied on rituals, organizational processes and tools in order to make consistent, inter-operable and reliable judgments (power_audit_1999), as well as provide broader access to a larger range of stakeholder participants in the audit process (birhane_power_2022). Audit tooling, then, represents not only a mechanism for maintenance and consistency of audit integrity but also a key capacity-building intervention to lower the barriers for broader participation.
Past work on AI audit tools. Recent research in human-computer interaction (HCI), social computing, and cooperative design documents the experiences of practitioners evaluating AI and the challenges they face (costanza-chock_who_2022; holstein_improving_2019; brown_toward_2019; madaio_co-designing_2020; devos_toward_2022; lee_landscape_2021). For example, holstein_improving_2019 documents the practical and technical difficulties faced by internal auditors of ML systems trying to identify and improve fairness, while other studies have examined the organizational challenges and barriers these practitioners face (rakova_where_2021; madaio_co-designing_2020; widder_its_2023; selbst_fairness_2019; costanza-chock_who_2022). Several HCI studies have since specifically examined the ways AI audit practitioners make use of various tools to address these issues (deng_exploring_2022; lee_landscape_2021; amershi_modeltracker_2015).
While some studies of AI audit tools focus on performance analysis (amershi_modeltracker_2015; harvey_gaps_2024) or user-driven grassroots auditing (devos_toward_2022; deng_understanding_2023), most HCI studies focus specifically on tools for assessing fairness and interpretability (bellamy_ai_2018; smith-renner_no_2020; kaur_interpreting_2020; bertrand_selective_2023; woodruff_qualitative_2018; madaio_co-designing_2020; lee_landscape_2021; deng_exploring_2022). lee_landscape_2021, for example, compare six prominent open source fairness toolkits along various criteria related to practitioners’ needs. deng_exploring_2022 documented the ways practitioners learn about and use two prominent fairness toolkits, AI Fairness 360 and Fairlearn (weerts_fairlearn_2024). And in a survey of 152 audit practitioners, costanza-chock_who_2022 found that 62% of practitioners used existing tools like AI Fairness 360, Scikit Fairness, or Parity, though only 7% of respondents used a standardized framework for their overall audit protocol.
However, less work examines tools for assessing harms beyond fairness, including refusal of medical services (waldman_how_2024), privacy violations, and other types of harm (shelby_sociotechnical_2023). And fairness toolkits do not typically support other necessary steps of an audit such as data collection. Our study aims to expand HCI research by looking beyond fairness evaluation toolkits to examine other kinds of tools involved in AI audits, such as tools for basic performance analysis or data collection.
We also build on more general critiques of “responsible AI” toolkits. wong_seeing_2023, for example, examine documentation from 27 toolkits for “AI Ethics”, finding that these resources employ a narrow technical framing that fails to involve more diverse stakeholders or reckon with the non-technical dimensions of AI ethics work; kaye_risky_2023 survey and critiques the tools currently used for AI governance by governments across the globe; and berman_scoping_2024 call for more evaluations of the effectiveness of responsible AI tools.
As yet, however, a comprehensive survey of auditors’ practical needs relative to the landscape of available tools is lacking. We extend these analyses and lay out a more complete view on what we refer to as AI audit tools: software, interfaces, code, benchmarks, frameworks, and other artifacts used by auditors in the AI audit process (Def. B.4). [^6] We include in this definition tools that may also be used for other “responsible AI” efforts, such as internal benchmarking, that do not meet our criteria for an audit (Def B.2). Auditing is an institutional arrangement—selecting the right tools does not guarantee operational independence, for example.
@@ 3. Methodology
To better understand the kinds of tools auditors use and where those tools fall short, we conducted 27 semi-structured interviews with 35 auditors across 24 organizations employing internal and external AI auditors. We use the term AI broadly to include tools applied to any AI-advertised product or model, including automated decision systems (ADS), algorithmic recommendation systems, large machine learning base models, generative AI products and more (see Fig. E.4). In parallel, to better understand existing tools, we curated a dataset of 435 tools designed or used for AI auditing and developed a taxonomy of the audit tool landscape based on our findings.
@@ 3.1. Interview methodology
We conducted 27 interviews with a total of 35 audit tool builders and practitioners, representing diverse backgrounds such as engineering, law, journalism, advocacy, policy, and academic research across North America ($N=22$) and Europe ($N=5$) (Table 1). These practitioners have all participated in internal or external audit work; many have also built tools for AI auditing. We used purposive sampling and snowball sampling methods to recruit participants. We began by contacting practitioners in our professional networks who had conducted notable AI audit work and were active in AI audit communities. Occasionally, participants were referred to us by a colleague or professional contact at another organization. Our sample encompassed both internal and external auditors employed by for-profit tech companies, AI evaluation startups, research and civil society non-profits, universities, and government agencies.
@@
Employer Roles of Interviewees Participants
Large tech for-profit Director of Policy Research, VP of Research, Data Science Mgr., Research Eng., Researcher P5, P8, P11, P14, P19, P27
Tech startup Co-Founder, CEO, Chief Scientist P4, P12
Government agency Tech Policy Principal/Mgr./Assoc./Advisor, Research Fellow (P21, P28-35), P25
University Assoc./Asst. Professor, Postdoc. Fellow, Data Scientist P3, P10, P13, P17, P20
Research non-profit Co-Founder, Director, Research Scientist P9, P16, P18
Civil society non-profit Director, Head of Analytics, Statistician, Researcher, Policy Fellow P1, P2, P6, P22, P24
Non-profit news org. Opinion Writer, Data Journalist P7, P15
Law/consulting for-profit Policy Director, Mgr., Consultant (P23, P28), P27
Interviews followed a semi-structured format and lasted 30–60 minutes. Our questions centered on 1) the specific tools and methods practitioners built or employed and 2) common obstacles and unmet needs. (The full interview protocol is included in Appendix D.) Participants had the option to remain anonymous and skip questions at their discretion, though none skipped a question. Our protocol was approved by three university IRBs. To analyze the interview data, we transcribed each interview and annotated the transcripts with manual codes. Our coding approach followed an inductive methodology, allowing patterns and themes to emerge from the data (wolcott_transforming_1994; glaser_discovery_2017). We employed a combination of descriptive coding, which captured the content of the interviews, and values coding, which captured the attitudes and beliefs expressed by participants. Through collaborative sessions and memo writing, we organized these codes and related quotes into key insights, presented in §4.